# Can Large Language Models Predict the Outcome of Judicial Decisions?

### v1

@@ VI Conclusion
In this study, we introduced the first Arabic instruction-following dataset for the LJP task and benchmarked LLMs on it. We evaluated approaches such as zero-shot, one-shot, and fine-tuning. Results show that smaller LLMs face challenges with this complex task, with minimal performance gains from prompt engineering. However, fine-tuning yielded significant improvements, demonstrating its effectiveness in adapting to the task. In-context learning (one-shot) also enhanced performance, achieving results close to fine-tuning, albeit slightly lower.
Future work could expand on this study by testing larger models to assess scaling effects, experimenting with advanced prompting methods, and exploring alternative fine-tuning strategies for further optimization. Additionally, the dataset provided here offers opportunities for the research community to develop new methodologies and applications in LJP for Arabic and other low-resource languages.

### v3

@@ VI Conclusion
This study introduced the first Arabic instruction-following dataset for LJP and benchmarked LLMs across different learning paradigms. Our results demonstrate that LoRA-based fine-tuning significantly enhances legal NLP performance, outperforming untrained and in-context approaches while reducing computational overhead. We showed the effectiveness of our approach for instruction generalization, to have a human-centric model.
Future work will explore the effectiveness of LLM-based scoring, larger models, and alternative fine-tuning methods. Additionally, developing a multi-domain, multilingual LLM for the legal domain could enhance model versatility and scalability across diverse legal contexts.